there are red and blue marbles in a bag. There are 34 more red marbles than there are blue. and there is a 108 total. how many blue are in the bag
Answer by lefty4ever26(59)

We will use the variables R for red and B for blue.
given information:
R=B+34
R+B=108 If we substitute R=B+34 into the second equation we can solve for how many blues there are.
(B+34)+B=108 Solve for B
2B+34=108
2B=74
B=37 Plug B back into the formula for R.

R=(37)+34
R=71
So, there are 37 blue marbles and 71 red marbles.
Hope this helps!